# Computing Richelot isogeny graph of superspecial abelian threefolds

### v1

@@ Abstract.
In algebraic geometry, superspecial curves are important research objects. While the number of superspecial genus-3 curves in characteristic $p$ is known, the number of hyperelliptic ones among them has not been determined even for small $p$. In this paper, in order to compute the latter number, we give an algorithm for computing the $(2,2,2)$-isogeny graph of superspecial abelian threefolds by using theta functions. This enables us to list up all superspecial genus-3 curves efficiently. Executing this algorithm on Magma, we succeeded in counting all superspecial hyperelliptic curves of genus $3$ when $10011\leq p<$ 100, and in finding a single one when $100007\leq p<$ 10000.
@@ 1. Introduction
Throughout this paper, by a curve we mean a non-singular projective variety of dimension one and isomorphisms between two curves are ones over an algebraically closed field. A curve $C$ over an algebraically closed field $k$ of characteristic $p>0$ is said to be superspecial if and only if the Jacobian of $C$ is isomorphic to the product of supersingular elliptic curves. Recently, superspecial curves are used for isogeny-based cryptography (e.g. [16], [5], [1] and so on), and then studies on such curves are becoming more important.
@@
[…]
Consequently, we obtain the following main theorem:
@@
Under some assumptions (see Remark 3.11 and Assumption 3.13), there is an algorithm (Algorithm 3.7) for computing the graph $\mathcal{G}_{3}(2,p)$ with $\widetilde{O}(p^{6})$ operations in the finite field $\hskip 0.85358pt\mathbb{F}_{p^{2}}$ for all chracteristics $p>7$.
@@
Executing our algorithm, we succeeded in solving the existence of superspecial hyperelliptic curves of genus 3 in small $p$ as follows:
@@
There exists a superspecial hyperelliptic curve of genus 3 in characteristic $p$ with $100007\leq p<$ 10000. In addition, the numbers of isomorphism classes of such curves for $10011\leq p<$ 100 are determined (as in Table 1 of Section 4.1).
@@
Acknowledgements
This research was conducted under a contract of “Research and development on new generation cryptography for secure wireless communication services” among “Research and Development for Expansion of Radio Wave Resources (JPJ000254)”, which was supported by the Ministry of Internal Affairs and Communications, Japan. This work was supported by JSPS Grant-in-Aid for Young Scientists 20K14301 and 23K12949, and WISE program (MEXT) at Kyushu University.
@@ 2. Preliminaries
In this section, we summarize background knowledge required in later sections. Abelian varieties and isogenies between them until Section 2.3 are defined over $\mathbb{C}$, but in Section 2.4 we consider those over a field of odd characteristic.
@@ 2.1. Abelian varieties and their isogenies
In this subsection, we review about complex abelian varieties and isogenies. A complex abelian variety  of dimension $g$ is isomorphic to $\mathbb{C}^{g}/(\mathbb{Z}^{g}+\varOmega\mathbb{Z}^{g})$ where $\varOmega$ is an element of the Siegel upper half-space
$\mathcal{H}_{g}\coloneqq\{\varOmega\in{\rm Mat}_{g}(\mathbb{C})\,\mid\,^{t}\varOmega=\varOmega,\,{\rm Im}\,\varOmega>0\}.\vspace{0.5mm}$
Then, we call $\varOmega\in\mathcal{H}_{g}$ a period matrix of the abelian variety. First of all, we recall a characterization of isomorphisms between abelian varieties:
@@
${\rm Sp}_{2g}(R)\coloneqq\{M\in{\rm GL}_{2g}(R)\,\mid\,ME\hskip 0.85358pt^{t}M=E\},\quad E\coloneqq\begin{pmatrix}0&{\mathbf{1}}_{g}\\ -{\mathbf{1}}_{g}&0\end{pmatrix}.$
@@
    Two matrices ${}^{t}\alpha\hskip 0.85358pt\gamma,\hskip 1.42262pt^{t}\beta\hskip 1.42262pt\delta$ are both symmetric and ${}^{t}\alpha\hskip 0.85358pt\delta-\hskip 1.42262pt^{t}\beta\hskip 1.42262pt\gamma={\mathbf{1}}_{g}$.
@@
$\mathbb{C}^{g}\longrightarrow\mathbb{C}^{g}\,;\,z\longmapsto^{t}(\gamma\varOmega+\delta)^{-1}z\eqqcolon M.z\vspace{0.5mm}$
@@
Theorem 2.3 ([28, Theorem 8.3.1]).
Two matrices $\varOmega,\varOmega^{\prime}\in\mathcal{H}_{g}$ determine isomorphic abelian varieties if and only if there exists $M\in{\rm Sp}_{2g}(\mathbb{Z})$ such that $\varOmega^{\prime}=M.\varOmega$.
An isogeny is a surjective homomorphism between two abelian varieties. If there exists an isogeny $A\rightarrow B$, then the dimensions of $A$ and $B$ are equal to each other, and we say that $A$ and $B$ are isogenous. Two isogenies $A\rightarrow B$ with the same kernel are equivalent up to an automorphism of $B$, and thus we identify them.
@@ 2.2. Theta functions
for $z\in\mathbb{C}^{g}$ and $\varOmega\in\mathcal{H}_{g}$. All the characteristics can be considered modulo $\mathbb{Z}^{g}$ since one can check that
$\theta\bigl{[}\begin{smallmatrix}a+\alpha\\ b+\beta\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)\exp(2\pi i\hskip 1.42262pt^{t}a\beta)$
for all $\alpha,\beta\in\mathbb{Z}^{g}$. Then, in the paper, we consider only theta functions $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)$ for characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$.
@@
The following two formulas are very fundamental and generate many relations:
@@
$\prod_{k=1}^{4}\theta[m_{k}](0,\varOmega)=\frac{1}{2^{g}}\sum_{\alpha,\beta\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}}(-1)^{4\hskip 0.56905pt^{t}a\beta}\prod_{k=1}^{4}\theta[n_{k}+\gamma_{\alpha,\beta}](0,\varOmega)\vspace{0.5mm}$
@@
Apply [23, Chapter IV, Theorem 1] for $z_{1}=z_{2}=z_{3}=z_{4}=0$. ∎
@@
Apply [23, Chapter IV, Theorem 2] for $m_{1}=m_{2}$ and $z_{1}=z_{2}=z$. ∎
@@
Lemma 2.8 ([9, Lemma F.3.1]).
With notations the above, we write $z=(z_{1},z_{2})$ where $z_{1}\in\mathbb{C}^{g_{1}},\,z_{2}\in\mathbb{C}^{g_{2}}$. Then, we have the equation
$\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}\\ b_{1}&b_{2}\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a_{1}\\ b_{1}\end{smallmatrix}\bigr{]}(z_{1},\varOmega_{1})\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}a_{2}\\ b_{2}\end{smallmatrix}\bigr{]}(z_{2},\varOmega_{2})\vspace{0.5mm}$
@@
$\theta\bigl{[}\begin{smallmatrix}M.a\\ M.b\end{smallmatrix}\bigr{]}(M.z,M.\varOmega)=c\cdot e^{\pi ik(M,a,b)}\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega),\vspace{1mm}$
@@
This follows from [28, Theorem 8.6.1]. ∎
@@ 2.3. Theta functions for abelian threefolds
In this subsection, we focus only on the case that $A$ is an abelian threefold (i.e.  abelian variety of dimension $g=3$), and we discuss theta functions associated to $A$. According to [28, Corollary 11.8.2], such $A$ is either of the following 4 types.
@@
$\vartheta_{i}(z,\varOmega):=\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ b_{1}&b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z,\varOmega),\quad i:=2b_{1}+4b_{2}+8b_{3}+16a_{1}+32a_{2}+64a_{3},\vspace{0.5mm}$
where $a_{1},a_{2},a_{3},b_{1},b_{2}$, and $b_{3}$ belong to $\{0,1/2\}$. Note that $i$ runs over $\{0,\ldots,63\}$. Moreover, the value $\vartheta_{i}(0,\varOmega)$ is simply be denoted as $\vartheta_{i}$.
We consider an abelian threefold $A\cong\mathbb{C}^{3}/(\mathbb{Z}^{3}+\varOmega\mathbb{Z}^{3})$, then the projective value
$[\vartheta_{0}^{2}:\vartheta_{1}^{2}:\cdots:\vartheta_{63}^{2}]\in\mathbb{P}^{63}(\mathbb{C})\vspace{0.5mm}$
is called a theta null-point of $A$. Remark that there are different theta null-points for isomorphic abelian threefolds.
@@
$S_{\rm van}\coloneqq\{\hskip 0.56905pti\in I_{\rm even}\mid\vartheta_{i}=0\}.$
@@
The value $N_{\rm van}$ is invariant for isomorphic abelian threefolds.
@@
This follows from the theta transformation formula (Theorem 2.9). ∎
Then, we can identify by $N_{\rm van}$ which of the above 4 types $A$ corresponds (remark that this does not depend on a period matrix $\varOmega$ of $A$ by the above lemma).
@@
[…]
@@
The first assertion follows from [18, Theorem 3.1] (the cases where $N_{\rm van}>9$ correspond to singular abelian threefolds). The second assertion for $N_{\rm van}=0,1$ are shown by [35, Theorem 2], and hence we show the remaining cases.
Let $A$ be an abelian threefold such that $A$ is isomorphic to the product $A_{1}\times A_{2}$ of an elliptic curve $A_{1}$ and an abelian surface $A_{2}$. Denote $\varOmega_{1},\varOmega_{2}$ and $\varOmega$ be one of period matrices of $A_{1},A_{2}$ and $A$, respectively. Then, there exists $M\in{\rm Sp}_{6}(\mathbb{Z})$ such that $M.\varOmega={\rm diag}(\varOmega_{1},\varOmega_{2})$ from Theorem 2.3, but we may assume $\varOmega={\rm diag}(\varOmega_{1},\varOmega_{2})$ by using Lemma 2.14. It follows from Lemma 2.8 that
$\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ b_{1}&b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a_{1}\\ b_{1}\end{smallmatrix}\bigr{]}(z_{1},\varOmega_{1})\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}a_{2}&a_{3}\\ b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z_{2},\varOmega_{2}),\quad z\coloneqq(z_{1},z_{2}),\vspace{0.5mm}$
and hence we see that $N_{\rm van}=6$ (resp.  $9$) if $A_{2}$ is simple (resp. not simple). ∎
At the rest of this subsection, we assume that an abelian threefold $A$ corresponds to a genus-3 curve $C$. Let us review how to restore a defining equation of $C$ in the case that $C$ is hyperelliptic.
@@
$C:y^{2}=x(x-1)(x-\lambda_{1})(x-\lambda_{2})(x-\lambda_{3})(x-\lambda_{4})(x-\lambda_{5}),$
@@
Apply Thomae’s formula (cf. [39]) for $g=3$ or see [36, Theorem 4]. ∎
On the other hand, even in the case that $C$ is a plane quartic, we can also restore an explicit equation of $C$. Taking a period matrix $\varOmega\in\mathcal{H}_{3}$ of $A={\rm Jac}(C)$, then it follows from Proposition 2.15 that $\vartheta_{i}\neq 0$ for all $i\in I_{\rm even}$. We define the following non-zero values
@@
computed as in [13, p. 15]. Then, there exist $k_{1},k_{2},k_{3},\lambda_{1},\lambda_{2},\lambda_{3}\in\mathbb{C}^{\times}$ such that
@@
Then, we can restore the curve $C$ from the values of $k_{i}$ using the following fact:
@@
[…]
2.4. Superspecial abelian threefold isogeny graph
Unlike the previous subsections, let us consider abelian varieties over a field of characteristic $p\geq 3$ in this subsection. An elliptic curve $E$ is called supersingular if $E[p]=\{0\}$.

### v2

# Listing superspecial curves of genus three by using Richelot isogeny graph

@@ Abstract.
In algebraic geometry, superspecial curves are important research objects. While the number of superspecial genus-3 curves in characteristic $p$ is known, the number of hyperelliptic ones among them has not been determined even for small $p$. In this paper, in order to compute the latter number, we give an algorithm for computing the Richelot isogeny graph of superspecial abelian threefolds by using theta functions. Our algorithm enables us to efficiently list superspecial genus-3 curves, and we succeeded in counting hyperelliptic curves among them when $10011\leq p<$ 100 by executing our algorithm in Magma.
@@ 1. Introduction
Throughout this paper, by a curve we mean a non-singular projective variety of dimension one and isomorphisms between two curves are ones over an algebraically closed field. A curve $C$ over an algebraically closed field $k$ of characteristic $p>0$ is called superspecial if the Jacobian of $C$ is isomorphic to the product of supersingular elliptic curves. In recent years, superspecial curves have been used for isogeny-based cryptography (e.g. [18], [6], [2] and so on), and studies on such curves are becoming more important.
@@
Our aim is to clarify the above problems for superspecial hyperelliptic curves of genus 3 in small characteristic $p$. For this, we utilize the connectivity of the Richelot isogeny graph (see Section 2.4 for definitions) of superspecial principally polarized abelian varieties of dimension $g$, denoted by $\mathcal{G}_{g}(2,p)$. While the graph $\mathcal{G}_{2}(2,p)$ can be computed by using known formulae (cf. [6, §$\S 3$]), this is not the case where $g\geq 3$ (partial formulae such as those in [50] exist for $g=3$, but they are not sufficient to compute the entire graph). In this paper, we will describe an explicit algorithm for computing $\mathcal{G}_{3}(2,p)$ by using theta functions. Theta functions enable us to compute Richelot isogenies between abelian varieties as outlined in [10, Appendix F], and the case of $g=2$ is implemented by [11] indeed. One of our contributions is explicitly formulating and implementing this algorithm in the case of $g=3$.
Our overall strategy for listing superspecial genus-3 curves is as follows: we start from one principally polarized superspecial abelian threefold, and compute vertices connected to it via edges on $\mathcal{G}_{3}(2,p)$, repeating this process to generate all vertices. An important point is that the computation of Richelot isogenies, the verification of whether the codomain is isomorphic to the Jacobian of a genus-3 curve, and the reconstruction of those curve can be carried out through theta constants. We note that this process is derived based on discussions over $\mathbb{C}$, but it also works correctly in characteristic $p>7$. Consequently, we obtain the following main theorem:
@@
There exists an algorithm (Algorithm 3.16) for listing superspecial genus-3 curves in characteristic $p>7$ with $\widetilde{O}(p^{6})$ operations over $\mathbb{F}_{p^{4}}$.
@@
Executing our algorithm, we succeeded in proving the existence of superspecial hyperelliptic curves of genus 3 in small $p$ as follows:
@@
The number of isomorphism classes of superspecial genus-3 hyperelliptic curves of genus 3 for $10011\leq p<$ 100 is summarized in Table 1 of Section 4. Moreover, there exists such a curve in any characteristic $p$ with $100007\leq p<$ 10000.
@@ 2. Preliminaries
In this section, we summarize background knowledge required in later sections. Abelian varieties and isogenies between them before Section 2.3 are defined over $\mathbb{C}$, but in Section 2.4 we consider those over a field of odd characteristic.
@@ 2.1. Abelian varieties and their isogenies
In this subsection, we shall review the theory of complex abelian varieties and isogenies. A complex abelian variety  of dimension $g$ is isomorphic to $\mathbb{C}^{g}/(\mathbb{Z}^{g}+\varOmega\mathbb{Z}^{g})$, where $\varOmega$ is an element of the Siegel upper half-space
$\mathcal{H}_{g}\coloneqq\{\varOmega\in{\rm Mat}_{g}(\mathbb{C})\,\mid\,^{t}\varOmega=\varOmega,\,{\rm Im}\,\varOmega>0\}.\vspace{1mm}$
Such $\varOmega\in\mathcal{H}_{g}$ is called a (small) period matrix of the abelian variety. First, we recall a characterization of isomorphisms between abelian varieties:
@@
${\rm Sp}_{2g}(R)\coloneqq\{M\in{\rm GL}_{2g}(R)\,\mid\,ME\hskip 0.85358pt^{t}M=E\},\quad E\coloneqq\begin{pmatrix}0&{\mathbf{1}}_{g}\\ -{\mathbf{1}}_{g}&0\end{pmatrix}.\vspace{0.7mm}$
@@
    Two matrices ${}^{t}\alpha\hskip 0.85358pt\gamma,\hskip 1.42262pt^{t}\beta\hskip 1.13809pt\delta$ are both symmetric and ${}^{t}\alpha\hskip 0.85358pt\delta-\hskip 0.28453pt^{t}\gamma\beta={\mathbf{1}}_{g}$.
@@
$\mathbb{C}^{g}\longrightarrow\mathbb{C}^{g}\,;\,z\longmapsto^{t}(\gamma\varOmega+\delta)^{-1}z\eqqcolon M.z\vspace{1mm}$
@@
Proposition 2.3 ([32, Proposition 8.1.3]).
The principally polarized abelian varieties of dimension $g$ with period matrices $\varOmega,\varOmega^{\prime}\in\mathcal{H}_{g}$ are isomorphic to each other if and only if there exists a symplectic matrix $M\in{\rm Sp}_{2g}(\mathbb{Z})$ such that $\varOmega^{\prime}=M.\varOmega$.
An isogeny between abelian varieties is a surjective homomorphism with a finite kernel. If there exists an isogeny $A\rightarrow B$, then the dimensions of $A$ and $B$ are equal to each other, and we say that $A$ and $B$ are isogenous. Two isogenies $A\rightarrow B$ with the same kernel are equivalent up to an automorphism of $B$, thus we identify them.
@@ 2.2. Theta functions
for $z\in\mathbb{C}^{g}$ and $\varOmega\in\mathcal{H}_{g}$. All the characteristics can be considered modulo $\mathbb{Z}^{g}$, since one can check that
$\theta\bigl{[}\begin{smallmatrix}a+\alpha\\ b+\beta\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)\exp(2\pi i\hskip 1.42262pt^{t}a\beta)\vspace{0.3mm}$
for all $\alpha,\beta\in\mathbb{Z}^{g}$. In the following, we shall consider only theta functions $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)$ for characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$.
@@
The following two formulae are very fundamental and generate many relations:
@@
$\prod_{k=1}^{4}\theta[m_{k}](0,\varOmega)=\frac{1}{2^{g}}\sum_{\alpha,\beta\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}}(-1)^{4\hskip 0.56905pt^{t}a\beta}\prod_{k=1}^{4}\theta[n_{k}+\gamma_{\alpha,\beta}](0,\varOmega)\vspace{0.7mm}$
@@
Apply [26, Chapter IV, Theorem 1] for $z_{1}=z_{2}=z_{3}=z_{4}=0$. ∎
@@
Apply [26, Chapter IV, Theorem 2] for $m_{1}=m_{2}$ and $z_{1}=z_{2}=z$. ∎
@@
Lemma 2.8 ([10, Lemma F.3.1]).
With the above notations, we write $z\coloneqq(z_{1},z_{2})$ where $z_{1}\in\mathbb{C}^{g_{1}},\,z_{2}\in\mathbb{C}^{g_{2}}$. Then, we have the equation
$\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}\\ b_{1}&b_{2}\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a_{1}\\ b_{1}\end{smallmatrix}\bigr{]}(z_{1},\varOmega_{1})\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}a_{2}\\ b_{2}\end{smallmatrix}\bigr{]}(z_{2},\varOmega_{2})\vspace{1mm}$
@@
$\theta\bigl{[}\begin{smallmatrix}M.a\\ M.b\end{smallmatrix}\bigr{]}(M.z,M.\varOmega)=c\cdot e^{\pi ik(M,a,b)}\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega),\vspace{0.7mm}$
@@
This follows from [32, Theorem 8.6.1]. ∎
@@ 2.3. Theta functions for abelian threefolds
In this subsection, we focus only on the case that $A$ is a principally polarized abelian threefold, and we discuss theta functions associated to $A$. Thanks to [32, Corollary 11.8.2], such $A$ is classified into one of the following four types:
@@
$\vartheta_{i}(z,\varOmega):=\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ b_{1}&b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z,\varOmega),\quad i:=2b_{1}+4b_{2}+8b_{3}+16a_{1}+32a_{2}+64a_{3},\vspace{0.8mm}$
where $a_{1},a_{2},a_{3},b_{1},b_{2}$, and $b_{3}$ belong to $\{0,1/2\}$. Note that $i$ runs over $\{0,\ldots,63\}$. For each $i$, the value $\vartheta_{i}(0,\varOmega)$ is written as $\vartheta_{i}$, and called the $i$-th theta constant.
We consider a principally polarized abelian threefold $A\cong\mathbb{C}^{3}/(\mathbb{Z}^{3}+\varOmega\mathbb{Z}^{3})$, then the projective value
$[\vartheta_{0}^{2}:\vartheta_{1}^{2}:\cdots:\vartheta_{63}^{2}]\in\mathbb{P}^{63}(\mathbb{C})\vspace{1mm}$
is called a (level 2) theta null-point of $A$. We remark that there are different theta null-points, even for isomorphic principally polarized abelian threefolds.
@@
$S_{\rm van}\coloneqq\{\hskip 0.56905pti\in I_{\rm even}\mid\vartheta_{i}=0\}.\vspace{0.3mm}$
@@
The value $N_{\rm van}$ in Definition 2.13 is invariant for isomorphic principally polarized abelian threefolds.
@@
This follows from Theorem 2.9 together with Proposition 2.3. ∎
Then, we can identify by $N_{\rm van}$ which of the above 4 types $A$ corresponds to (note that this does not depend on a period matrix $\varOmega$ of $A$ by the above lemma).
@@
$\left\{\begin{array}[]{l}N_{\rm van}=0\ \Leftrightarrow\,A\cong{\rm Jac}(C)\ \text{where}\ C\ \text{is\ a\ smooth\ plane\ quartic},\\[1.42262pt] N_{\rm van}=1\ \Leftrightarrow\,A\cong{\rm Jac}(C)\ \text{where}\ C\ \text{is\ a\ hyperelliptic\ curve\ of\ genus\ }3,\\[1.42262pt] N_{\rm van}=6\ \Leftrightarrow\,A\cong E\times{\rm Jac}(C)\ \text{where}\ E\ \text{and}\ C\text{\ is\ a genus-1 and genus-2 curve},\\[1.42262pt] N_{\rm van}=9\ \Leftrightarrow\,A\cong E_{1}\times E_{2}\times E_{3}\ \text{where}\ E_{1},E_{2}\text{\ and\ }E_{3}\text{\ are\ elliptic\ curves}.\end{array}\right.\vspace{-0.3mm}$
@@
The first assertion follows from [20, Theorem 3.1] (the cases where $N_{\rm van}>9$ correspond to singular abelian threefolds). Also, for the second assertion, the cases where $N_{\rm van}=0,1$ follows from [42, Theorem 2]. Note that these results are proved over $\mathbb{C}$, but valid over an algebraically closed field of characteristic $\neq 2$ as stated in Pieper’s paper [43, Remark 6.3]. In the following, we show the remaining cases.
Let $A$ be a principally polarized abelian threefold isomorphic to the (principally polarized) product of an elliptic curve $A_{1}$ and an abelian surface $A_{2}$. Denote $\varOmega_{1},\varOmega_{2}$, and $\varOmega$ to be the period matrices of $A_{1},A_{2}$, and $A$ respectively. Theorem 2.3 shows that there exists $M\in{\rm Sp}_{6}(\mathbb{Z})$ such that $M.\varOmega={\rm diag}(\varOmega_{1},\varOmega_{2})$, hence we may assume that $\varOmega={\rm diag}(\varOmega_{1},\varOmega_{2})$ using Lemma 2.14. Then, it follows from Lemma 2.8 that
$\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ b_{1}&b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a_{1}\\ b_{1}\end{smallmatrix}\bigr{]}(z_{1},\varOmega_{1})\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}a_{2}&a_{3}\\ b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z_{2},\varOmega_{2}),\quad z\coloneqq(z_{1},z_{2}).\vspace{1mm}$
As known in [34, Remark 3.5] or [46, §$\S 16.4$], a simple (resp. not simple) principally polarized abelian surface has no (resp. exactly one) vanishing even theta constant. This shows that $N_{\rm van}=6$ if $A_{2}$ is simple and $N_{\rm van}=9$ otherwise, as desired. ∎
In the rest of this subsection, suppose that $A$ corresponds to a genus-3 curve $C$. Let us review how to restore a defining equation of $C$ from theta constants of $A$.
@@
$C:y^{2}=x(x-1)(x-\lambda_{1})(x-\lambda_{2})(x-\lambda_{3})(x-\lambda_{4})(x-\lambda_{5}),\vspace{-0.2mm}$
@@
Apply [51, Theorem 1.1] for $g=3$, or see [47, Theorem 4]. ∎
On the other hand, even in the case that $C$ is a plane quartic, we can also restore a defining equation of $C$ as follows: Taking a period matrix $\varOmega\in\mathcal{H}_{3}$ of $A={\rm Jac}(C)$, it follows from Proposition 2.15 that $\vartheta_{i}\neq 0$ holds for all $i\in I_{\rm even}$. Then, we define the following non-zero values
@@
computed as in [15, p. 15]. Then, there exist $k_{1},k_{2},k_{3},\lambda_{1},\lambda_{2},\lambda_{3}\in\mathbb{C}^{\times}$ such that
@@
See [54, §$\S 15$], [39, Theorem 3.1], or [15, Proposition 2]. ∎
2.4. Superspecial Richelot isogeny graph
Unlike the previous subsections, let us consider abelian varieties over an algebraically closed field of characteristic $p\geq 3$ in this subsection. An elliptic curve $E$ is called supersingular if $E[p]=\{0\}$.